A coin bank contains $17.00 in pennies and nickels. If there are 1140 coins in the bank how many of the coins are nickels?
Answer by CharlesG2(834)

A coin bank contains $17.00 in pennies and nickels. If there are 1140 coins in the bank how many of the coins are nickels?
P pennies, N nickels
P + N = 1140
0.01P + 0.05N = 17 --> P + 5N = 1700 (multiplied both sides by 100)
P = 1700 - 5N
1700 - 5N + N = 1140
1700 - 4N = 1140
560 = 4N
140 nickels = N
1140 - 140 = 1000 = P
0.01 * 1000 = 10
0.05 * 140 = 7
10 + 7 = 17